4. Sam has 6 rose bushes. He counted the flowers on each of them. There are 82,5 4, 11 and 9. Find the Standard Deviation. X a "Usualnumber of flowers? X is your last two digits of your CUNYFirst ID Number. (00=0,01=1,...)(20 points)


1
Expert's answer
2021-06-14T14:56:58-0400
"8, 2, 5, 4, 11, 9"


"mean=\bar{x}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nx_i"

"=\dfrac{1}{6}(8+2+5+4+11+9)=6.5"


"Var(X)=\sigma^2=\dfrac{1}{n}\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2"

"=\dfrac{1}{6}((8-6.5)^2+(2-6.5)^2+(5-6.5)^2+(4-6.5)^2"

"+(11-6.5)^2+(9-6.5)^2)=\dfrac{57.5}{6}\approx9.583333"

"\sigma=\sqrt{\sigma^2}\approx3.0957"